OpenStudy (anonymous):

Suppose y varies directly with x and y=72when x=6 . What direct variation equation relates x and y? What is the value of y when x=10

OpenStudy (anonymous):

y = k*x 72 = k*6 k=72/6= 12 y = 12*x when x = 10 what is y?

OpenStudy (anonymous):

@loppol do u kno?

OpenStudy (anonymous):

its doesnt say

OpenStudy (anonymous):

i think 120 if 10 times 12

OpenStudy (anonymous):

corret 12*10=120